A buffer that contains 1.05 M base, B, and 0.750 M of its conjugate acid, BH+, has a pH of 9.250. What is the pH after 0.0055 mol of HCl is added to 0.400 L of this solution?
Answer #1

Before [B]=1.05M [BH+ = 0.75M pH = 9.25 polt=144 - pH = 14-9-25=4.75 polt=pkb tog [B] 0.75 4.75= plb + logo 0.15 :4.826 -os P